OpenStudy (anonymous):

For f(x)=2x+1 and g(x)=x^2-7, find (f-g)(X) a:-x^2+2x-6 b:2x^2-15 c:x^2-2x-8 d:8+2x-x^2

OpenStudy (anonymous):

subtract both functions

OpenStudy (anonymous):

Can you set that up for me? I know you have to subtract them, although I have no idea how to set it up

OpenStudy (welshfella):

f(x) - g(x) = 2x+ 1 - (x^2 - 7)

OpenStudy (welshfella):

take care when you distribute the - over the parentheses

OpenStudy (welshfella):

what do you get?

OpenStudy (anonymous):

i accidentally clicked the wrong answer

OpenStudy (jhannybean):

Think about it like this: let a = 2x+1 , b = x\(^2\)-7 . So you're essentially looking for a-b.

OpenStudy (welshfella):

here's an example - (x - 3) = -1*x -1*-3 = - x + 3 - just as if you were distributing -1